OpenStudy (anonymous):

pleaseeee help What is 3 3/8 x 4 1/3 ? Express the answer in simplest form.

jabez177 (jabez177):

First we have to convert the mixed fraction to an improper fraction.

jabez177 (jabez177):

3 x 8 + 3 is?

OpenStudy (anonymous):

27?

jabez177 (jabez177):

Good, now the other mixed fraction. 4 x 3 + 1 = ?

OpenStudy (anonymous):

13?

jabez177 (jabez177):

\[\frac{ 27 }{ 8 }\times \frac{ 13 }{ 3 }\]

jabez177 (jabez177):

3 goes into 27 how many times? Meaning, what times 3 equals 27?

OpenStudy (anonymous):

9 times?

jabez177 (jabez177):

Good, now cross off the 27 and put 9 there and where the 3 is, put a 1. Like this, \[\frac{ 9 }{ 8 }\times \frac{ 13 }{ 1 }\]

jabez177 (jabez177):

Now, 9 x 13 is?

OpenStudy (anonymous):

117?

jabez177 (jabez177):

Good now what is 8 x 1?

OpenStudy (anonymous):

8

jabez177 (jabez177):

Okay, now we have to divide 117 by 8.

jabez177 (jabez177):